# Large-scale, Independent and Comprehensive study of the power of LLMs for test case generation

### v1

@@ 3.5. The EvoSuite SBST tool
To account for the inherent randomness of genetic algorithms used by EvoSuite, we executed the tool 30 times for each class under test. This aligns with previous research suggesting a three-minute time budget per test case generation run (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015). This time limit reflects practical considerations faced by developers during the testing process.
@@ 4. Experimental Results
∗Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for the small dataset (CMD).
@@
[Experiment Design]: We consider five prompt designs to instruct the LLMs to generate test cases and use regular expressions to extract the source code of test suites from the LLM outputs. To evaluate the effectiveness of the techniques, we adopted the Match Success Rate (MSR) and Code Extraction Success Rate (CSR) metrics introduced by Macedo et al. (Macedo et al., 2024). MSR measures the proportion of outputs where the output could be matched by a pre-defined regexp pattern. It is thus a proxy to measure to what extent the LLM adheres to the format specified in the prompt. CSR computes the proportion of cases where a source code can be extracted within the matched output.
@@
[Experiment Results]: Our study generated 216 300 files, with 190 619 containing source code and 171 187 cases where syntactically correct test suites are provided. Table 3 provides performance measurements of LLMs in terms of MSR and CSR.
@@
Table 4 summarizes the performance of LLMs in generating syntactically correct (i.e., parsable) and compilable test suites. Overall, the metrics clearly show that LLMs underperform in generating syntactically correct test cases and are further struggling to generate compilable ones. In the best case scenario, GPT-3.5-Turbo with GToT prompting offers the highest effectiveness score on the Defects4J dataset with only 24.54% of generated test cases being syntactically correct, and 6.76% of these being also compilable. Overall, the results show that, on average, only 12% of the syntactically correct test suites compile successfully.
Following compilation, we perform static analysis using SpotBugs, which identifies 8 461 potential issues across 20 530 test cases. We classify these issues (aka bugs) into four severity levels: Scariest (most critical, affecting functionality and security), Scary (significant issues if unaddressed), Troubling (minor issues), and Of Concern (indicating lower severity). The majority (55.58%) of the issues are labeled as only Of Concern, i.e., they are informational and of minimal risk. We find that Few-shot prompts lead the LLMs to generate most (61.72%) of the ’Scariest’ bugs. We nevertheless note that GToT is associated with most of the ’Scary’ (63.92%), ’Troubling’ (69.10%), and ’Of Concern’ (33.77%) bugs. Table 5 details the statistics of issues found by SpotBugs in the test suites generated by GPT-3.5-turbo using varying prompts. ToT generates the least issues while CoT generates only ’of concern’ issues. In contrast, FSL leads to a large number of test cases with ’Scariest’ issues.
| Rank-Level | ZSL | FSL | CoT | ToT | GToT | Total |
| Of Concern | 844 | 619 | 794 | 51 | 1106 | 3414 |
| Scariest | 0 | 587 | 0 | 1 | 0 | 588 |
| Scary | 0 | 0 | 0 | 2 | 825 | 827 |
| Troubling | 227 | 14 | 0 | 3 | 825 | 1069 |
@@
– Sun Code Conventions: For Sun coding standards, Few-shot learning was again the most compliant, followed by CoT and Zero-shot learning. ToT and GToT had higher violations, indicating weaker adherence and readability. Common issues (Figure 4(b)) included inadequate method documentation, use of ”magic” numbers, regex problems, excessive line length, and insufficient package documentation. These findings highlight the impact on code maintainability and clarity.
@@
| Cyclomatic Complexity Level | ZSL | FSL | CoT | ToT | GToT |
| Low complexity (1-4) | 3 601 | 2 794 | 3 669 | 2 960 | 7 468 |
| Moderate complexity (5-7) | 4 | 11 | 6 | 6 | 1 |
| High complexity (8-10) | 1 | 3 | 4 | 2 | 0 |
| Very High complexity (11+) | 0 | 0 | 0 | 0 | 0 |
| Cognitive Complexity Level | ZSL | FSL | CoT | ToT | GToT |
| Low complexity (¡5) | 3 600 | 2 804 | 3 674 | 2 967 | 7 469 |
| Moderate complexity (5-10) | 4 | 4 | 5 | 1 | 0 |
| High complexity (11-20) | 1 | 0 | 0 | 0 | 0 |
| Very High complexity (21+) | 1 | 0 | 0 | 0 | 0 |
@@
– Cognitive Complexity: Most test suites generated exhibit low complexity (Table 7(b)), particularly those produced by the Guided Tree-of-thoughts (GToT) technique, which produced the most low complexity cases. This demonstrates GToT’s ability to generate simple and maintainable code. Instances of moderate to very high complexity are rare, indicating the overall effectiveness of these techniques in producing straightforward and understandable code.
– Cyclomatic Complexity: Most generated test suites, especially with the GToT technique, exhibit low complexity (Table 7(a)), indicating effective, maintainable code creation. Moderate and high complexity instances are present but don’t reach very high levels, ensuring good maintainability. Other techniques also produce manageable code but may occasionally result in higher complexity, impacting maintainability and comprehension.
@@
LLMs exhibit variable performance across different metrics, particularly GPT-3.5 Turbo, which is heavily influenced by advanced prompting techniques like GToT and CoT (Figure 5). In the Custom dataset, GPT-3.5 Turbo achieves near-perfect median coverage (100%), suggesting effective adaptation to novel software contexts. Interestingly, it outperforms GPT-4 in this dataset, potentially due to the smaller size or limitations of GPT-4 for this specific task. However, GPT-3.5 Turbo’s performance varies across other datasets. It demonstrates a low median line coverage of 3.57% in SF110 but achieves significantly higher coverage in Defects4J (82.74% lines, 85.0% instructions, 100% methods) compared to SF110 (2.0% lines, 16.67% instructions). Mistral 7B also shows inconsistent performance, particularly struggling in SF110 with low median coverages (3.03% lines, 2.0% instructions, 20.0% methods). However, it achieves improved results in Defects4J, with higher median coverages (63.4% lines, 66.0% instructions, 100% methods).
@@
Our findings highlight the importance of refining prompt engineering to fully leverage LLM capabilities across diverse testing scenarios. Improved prompting, combined with exploring SBST methodologies, shows potential for enhancing test suite quality and effectiveness. Additionally, the superior performance of GPT-3.5 Turbo over GPT-4 in the Custom dataset requires cautious interpretation. Further analysis is needed to determine if this is due to actual model advancements or the limited sample size.
4.4. [RQ4]: Bug Detection.
@@
[Experiment Goal]: Our aim is to evaluate to what extent the test suites generated by LLMs can detect bugs compared with those produced by an SBST method.
[Experiment Design]:
We employed two controlled experiments to assess the bug detection capabilities of the targeted models. The first experiment leveraged the Defects4J benchmark (Just et al., 2014), a well-established dataset encompassing 403 real-world software bugs across diverse Java projects. This experiment compared the performance of three models: GPT-3.5 turbo, Mistral 7B, and EvoSuite, a leading SBST tool. We rely only on Defects4J since SF110 does not provide ground truth fixed versions of programs to validate bug detection.
The second experiment builds on a Custom dataset we developed, containing 20 real-world software bugs extracted from recent (post-May 2023) GitHub projects. This dataset was necessary to evaluate the generalizability of the findings on unfamiliar datasets for the LLMs. The second experiment included four models: GPT-3.5 turbo, GPT-4 (a more advanced version of GPT-3.5), Mistral 7B, and Mixtral 8x7B (a potentially higher-capacity variant of Mistral 7B that we use sparingly due to budget constraints).
The evaluation procedure involved running the test suite on the buggy version of the class. If the test suite passes, the bug is undetected; if it fails, it is run on the fixed version. A passing result on the fixed version confirms bug detection. If it still fails, error messages and the number of failing tests are analyzed. A decrease in failing tests on the fixed class indicates successful bug detection; otherwise, the bug remains undetected.
[Experiment Results]: Table 8 presents the results of bug detection with the test suites generated by Mistral 7B, GPT 3.5-turbo and EvoSuite on the Defects4J dataset. GPT 3.5-turbo and Mistral 7B showed varied effectiveness across different prompting techniques, with CoT achieving the highest bug detection rate. Specifically, CoT with Mistral 7B detected 150 out of 403 bugs, outperforming other techniques. In contrast, EvoSuite detected only 3 out of 403 bugs. This discrepancy raises questions about the conventional applications and configurations of SBST tools like EvoSuite, which are primarily designed for regression testing scenarios. These tools may not be optimized for detecting bugs in a forward testing scenario, as evidenced by the low detection rates observed.
On the Custom dataset, all LLMs failed to detect bugs. This failure suggests that either the Defects4J dataset, in contrast to CMD, includes easy-to-detect bugs or the test cases of Defects4J have been memorized by the LLMs. LLMs also likely falter when prompted on novel software frameworks or languages that introduce less common syntax or constructs. Despite many classes in the custom dataset being of low complexity (Table 2), moderately and highly complex classes suggest that inherent complexities might further challenge effective bug detection. Finally, the high bug detection rates with ZSL prompts, contrary to FSL and even ToT, also cast suspicion of data leakage. CoT prompts exhibit the strongest overall performance regarding bug detection capabilities of the generated test suites. This is likely due to CoT’s effective simulation of real-world testing scenarios, which encourages deeper exploration of code logical flows to uncover bugs.
| Projects | # bugs | Mistral 7B | GPT 3.5-turbo | EvoSuite | ||||||||
| CoT | ToT | GToT | FSL | ZSL | CoT | ToT | GToT | FSL | ZSL | SBST | ||
| Cli | 30 | 17 | 14 | 14 | 14 | 16 | 20 | 19 | 14 | 18 | 20 | 1 |
| Closure | 63 | 22 | 11 | 10 | 10 | 17 | 18 | 8 | 11 | 5 | 12 | 0 |
| Codec | 8 | 3 | 2 | 2 | 2 | 2 | 3 | 2 | 2 | 2 | 2 | 0 |
| Collections | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 0 |
| Compress | 22 | 10 | 9 | 9 | 8 | 10 | 10 | 10 | 11 | 9 | 12 | 1 |
| Csv | 9 | 5 | 5 | 5 | 5 | 6 | 6 | 4 | 5 | 4 | 6 | 0 |
| Gson | 9 | 4 | 0 | 0 | 0 | 3 | 4 | 0 | 0 | 0 | 2 | 0 |
| JacksonCore | 9 | 0 | 1 | 0 | 2 | 1 | 1 | 2 | 2 | 2 | 2 | 0 |
| JacksonDatabind | 68 | 1 | 0 | 0 | 1 | 1 | 0 | 0 | 0 | 0 | 1 | 0 |
| JacksonXml | 2 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Jsoup | 61 | 41 | 28 | 28 | 26 | 33 | 41 | 26 | 26 | 26 | 28 | 0 |
| JxPath | 14 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| Lang | 14 | 12 | 11 | 11 | 11 | 12 | 11 | 10 | 10 | 10 | 11 | 0 |
| Math | 54 | 15 | 6 | 6 | 7 | 11 | 9 | 5 | 6 | 5 | 7 | 0 |
| Mockito | 33 | 19 | 5 | 5 | 6 | 9 | 19 | 5 | 4 | 6 | 6 | 0 |
| Time | 6 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 0 | 0 | 0 | 1 |
| Total | 403 | 150 | 93 | 91 | 93 | 128 | 143 | 93 | 92 | 88 | 110 | 03 |
@@
Models include Mistral 7B, GPT 3.5-turbo, and EvoSuite. ZSL: Zero-shot learning, FSL: Few-shot learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts, SBST: Search-Based Software Testing.

### v2

@@ 3.5. The EvoSuite SBST tool
To account for the inherent randomness of genetic algorithms used by EvoSuite, we executed the tool 30 times for each class under test like in (Tang et al., 2024). We allocated a three-minute time budget for each test case generation, chosen to reflect practical constraints typically encountered by developers. Previous studies (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015) have shown that a three-minute time budget is generally sufficient for effective test case generation.
@@ 3.6. Datasets
ODC, with 13 releases, 527 stars, and 20 contributors, is a collaborative database development tool offering features such as SQL development, risk management, and data security. Conductor OSS, originally developed by Netflix, is a microservices orchestration platform with 11 releases, 15.5k stars, and 53 contributors. These projects were selected due to their complexity and active development, providing a robust basis for testing LLM-generated unit tests. CMD was limited to two projects to accommodate the token limits of the LLMs, but future work will expand the dataset to improve generalizability.
@@ 4. Experimental Results
∗Note that due to budget constraints, we only paid to query GPT-4 and Mixtral 8x7B for CMD.
@@
[Experiment Design]: We consider five prompt designs to instruct the LLMs to generate test cases and use regular expressions to extract the source code of test suites from the LLM outputs. To evaluate the effectiveness of the techniques, we adopted the Match Success Rate (MSR) and Code Extraction Success Rate (CSR) metrics introduced by Macedo et al. (Macedo et al., 2024). MSR measures the proportion of outputs where the output could be matched by a pre-defined regexp pattern. It is thus a proxy to measure to what extent the LLM adheres to the format specified in the prompt. CSR computes the proportion of cases where a source code can be extracted within the matched output. We opted to use regular expressions as a lightweight and efficient solution. This approach is consistent with prior work (Macedo et al., 2024).
@@
[Experiment Results]: Our study generated a total of 216 300 files, of which 190 619 contained source code. Among these, 171 187 were syntactically correct test suites, while 25 681 files contained only textual outputs or incorrect data. Table 3 provides performance measurements of LLMs in terms of MSR and CSR.
@@
Table 4 summarizes the performance of LLMs in generating syntactically correct (i.e., parsable) and compilable test suites. Overall, the metrics clearly show that LLMs underperform in generating syntactically correct test cases and are further struggling to generate compilable ones. In the best case scenario, GPT-3.5-Turbo with GToT prompting offers the highest effectiveness score on the Defects4J dataset with only 24.54% of generated test cases being syntactically correct, and 6.76% of these being also compilable. Overall, the results show that, on average, only 12% of the syntactically correct test suites compile successfully. The discrepancy between the total number of generated files and the number of syntactically correct test suites stems from the complexity of generating tests at the class level. Class-level generation, which involves generating tests for entire classes rather than individual methods, introduces additional challenges, including higher susceptibility to compilation errors. As a result, not all generated files contained valid or compilable test suites, contributing to the difference between the total generated files and the correctly compiled tests.
Following compilation, we perform static analysis using SpotBugs, which identifies 8 461 potential issues across 20 530 test cases (Table 5). We classify these issues (aka bugs) into four severity levels: Scariest (most critical, affecting functionality and security), Scary (significant issues if unaddressed), Troubling (minor issues), and Of Concern (indicating lower severity). The majority (55.58%) of the issues are labeled as only Of Concern, i.e., they are informational and of minimal risk. We find that Few-shot prompts lead the LLMs to generate most (61.72%) of the ’Scariest’ bugs. We nevertheless note that GToT is associated with most of the ’Scary’ (63.92%), ’Troubling’ (69.10%), and ’Of Concern’ (33.77%) bugs.
@@
– Sun Code Conventions: For Sun coding standards, Few-shot learning is again the most compliant, followed by CoT and Zero-shot learning. ToT and GToT had higher violations, indicating weaker adherence and readability. Common issues (Figure 4(b)) included inadequate method documentation, use of ”magic” numbers, regex problems, excessive line length, and insufficient package documentation. These findings highlight the impact on code maintainability and clarity.
@@
| Cyclomatic Complexity | ZSL | FSL | CoT | ToT | GToT |
| Low (1-4) | 3 601 | 2 794 | 3 669 | 2 960 | 7 468 |
| Moderate (5-7) | 4 | 11 | 6 | 6 | 1 |
| High (8-10) | 1 | 3 | 4 | 2 | 0 |
| Very High (11+) | 0 | 0 | 0 | 0 | 0 |
| Cognitive Complexity | ZSL | FSL | CoT | ToT | GToT |
| Low (¡5) | 3 600 | 2 804 | 3 674 | 2 967 | 7 469 |
| Moderate (5-10) | 4 | 4 | 5 | 1 | 0 |
| High (11-20) | 1 | 0 | 0 | 0 | 0 |
| Very High (21+) | 1 | 0 | 0 | 0 | 0 |
@@
– Cognitive Complexity: Most test suites generated exhibit low complexity, particularly those produced by the Guided Tree-of-Thoughts (GToT) technique, which resulted in 7,469 low-complexity cases (Table 6(b)), the highest among all techniques. This accounts for over 99% of the tests generated by GToT, demonstrating its ability to consistently generate simple and maintainable code. Instances of moderate to very high complexity are rare, with only 1 case in the moderate range and none in the high or very high ranges, further highlighting the overall effectiveness of these techniques in producing straightforward and understandable code.
– Cyclomatic Complexity: Most generated test suites also exhibit low cyclomatic complexity, especially with the GToT technique, which produced 7,468 low-complexity cases (Table 6(a)), accounting for nearly all the test suites generated by this technique. Moderate and high complexity instances are minimal, with just 1 moderate-complexity case and no high-complexity cases. Other techniques like CoT and ZSL also result in a majority of low-complexity test suites, though they have slightly higher instances of moderate complexity (6 and 4 cases, respectively). This indicates that while most techniques generate maintainable code, GToT stands out for producing simpler, more maintainable test suites.
@@
LLMs exhibit variable performance across different metrics, particularly GPT-3.5 Turbo, which is heavily influenced by advanced prompting techniques like GToT and CoT (Figure 5). In the Custom dataset, GPT-3.5 Turbo achieves near-perfect median coverage (100%), suggesting effective adaptation to novel software contexts. Interestingly, GPT-3.5 Turbo performs better than GPT-4 on this dataset, potentially due to the smaller size or specific limitations of GPT-4 for this particular task. However, GPT-3.5 Turbo’s performance varies across other datasets. It demonstrates a low median line coverage of 3.57% in SF110 but achieves significantly higher coverage in Defects4J (82.74% lines, 85.0% instructions, 100% methods) compared to SF110 (2.0% lines, 16.67% instructions). Mistral 7B also shows inconsistent performance, particularly struggling in SF110 with low median coverages (3.03% lines, 2.0% instructions, 20.0% methods). However, it achieves improved results in Defects4J, with higher median coverages (63.4% lines, 66.0% instructions, 100% methods).
@@
Our findings highlight the importance of refining prompt engineering to fully leverage LLM capabilities across diverse testing scenarios. Improved prompting, combined with exploring SBST methodologies, shows potential for enhancing test suite quality and effectiveness. Additionally, the better performance of GPT-3.5 Turbo over GPT-4 in the Custom dataset requires cautious interpretation. Further analysis is needed to determine if this is due to actual model advancements or the limited sample size.
4.4. [RQ4]: Test Smell detection.
@@
[Experiment Goal]: Our aim is to evaluate the extent to which the test suites generated by LLMs exhibit test smells compared with those produced by an SBST method like EvoSuite.
[Experiment Design]: This experiment assesses the test smell prevalence in test suites generated by LLMs and EvoSuite. CMD was introduced to evaluate the generalizability of LLM-generated test suites on modern and unfamiliar codebases. GPT-4 and Mixtral-8x7B were evaluated only on the CMD due to budget constraints and the focus on understanding the models’ performance on novel data. For EvoSuite, the experiment was applied only to the SF110 and Defects4J, as EvoSuite is incompatible with CMD (cf. Section 3.6). Test smells were detected using TsDetect (Peruma et al., 2020), a state-of-the-art tool that identifies 20 different types of test smells (Peruma et al., 2019, 2020). Due to space constraints, we provide a list of the test smells detectable by TsDetect with their descriptions in our replication package.
The evaluation involved analyzing the test suites for common test smells, such as incomplete assertions, complex test logic, or excessive setup methods. We recorded the frequency and types of test smells detected in each test suite, allowing us to compare the absence of test smells of LLM-generated tests against those produced by EvoSuite.
[Experiment Results]: Table 7 shows that the most common test smells (Van Deursen et al., 2001; Meszaros et al., 2003; Peruma et al., 2019, 2020) produced by both LLMs and EvoSuite are Assertion Roulette (AR), Conditional Logic Test (CLT), Constructor Initialization (CI), Empty Test (EM), Exception Handling (EH), General Fixture (GF), Mystery Guest (MG), and Magic Number Test (MNT), among others. Smells like Default Test, Verbose Test, and Dependent Test were not found in any tests. The results reveal notable differences in test quality across prompt techniques and models, with techniques like COT and GToT producing higher-quality tests, while Few-shot Learning and EvoSuite tend to generate tests with more test smells. The values marked as ”-” for Mistral 7B CMD with ZSL and CoT in Table 7 are due to the absence of compilable test cases for these particular prompt techniques.
For GPT-3.5 Turbo, CoT produces relatively clean tests, with Assertion Roulette in 40.06% of Defects4J files, though Magic Number Test remains high at 99.28%. Few-shot Learning performs worse, with 73.15% affected by Assertion Roulette. On CMD, CoT reduces Assertion Roulette to 6.4%, but Magic Number Test persists at 100%. GToT further reduces Assertion Roulette to 19.86%, though Magic Number Test remains. With GPT-4, the trends are similar, with CoT reducing Assertion Roulette to 14.56% on CMD but Magic Number Test still at 98%. ToT and GToT show slightly lower Assertion Roulette but struggle with Magic Number Test.
Regarding Mistral7B, CoT and Few-shot Learning result in high Assertion Roulette on Defects4J (65.13% and 79.35%), while GToT reduces it to 27.97%. On CMD, GToT performs better, reducing Assertion Roulette to 16.67%, though Magic Number Test persists in 66.67% of files. For Mixtral8x7b, used on CMD, CoT generates few smells, but Magic Number Test remains at 100%. Few-shot Learning fares better with 33.33% showing Assertion Roulette, but Magic Number Test persists. Zero-shot Learning is less effective, with 75% showing Assertion Roulette.
EvoSuite shows similar trends, with 38.24% Assertion Roulette on Defects4J and 100% Magic Number Test. Other significant smells include General Fixture (76.94%) and Redundant Assertion (54.67%) on Defects4J, and higher Assertion Roulette (50.78%) on SF110, along with General Fixture (62.95%) and Lazy Test (6.73%).
| Model | Dataset | Prompt | Test Smell (%) | |||||||||||||||||
| AR | CLT | CI | EM | EH | GF | MG | RP | RA | SE | ST | EA | LT | DA | UT | IT | RO | MNT | |||
| GPT 3 | SF110 | ZSL | 46.17 | 2.68 | 0.16 | 12.71 | 13.02 | 11.13 | 2.92 | 0.0 | 0.79 | 23.52 | 0.39 | 48.38 | 53.75 | 9.31 | 27.70 | 0.0 | 3.24 | 99.84 |
| FSL | 74.44 | 3.91 | 0.11 | 7.14 | 14.51 | 6.70 | 3.57 | 0.0 | 3.13 | 33.15 | 0.45 | 70.65 | 53.91 | 16.52 | 16.29 | 0.0 | 4.13 | 100 | ||
| CoT | 40.98 | 2.15 | 0.0 | 22.25 | 11.60 | 5.93 | 1.98 | 0.0 | 0.95 | 17.53 | 0.09 | 42.44 | 50.26 | 6.53 | 34.54 | 0.0 | 1.89 | 99.05 | ||
| ToT | 59.77 | 3.22 | 0.11 | 4.94 | 19.20 | 13.45 | 2.41 | 0.34 | 3.56 | 28.51 | 0.92 | 56.90 | 76.21 | 13.33 | 15.06 | 0.11 | 2.64 | 100 | ||
| GToT | 38.23 | 0.58 | 0.0 | 28.97 | 10.24 | 2.77 | 1.99 | 0.0 | 2.13 | 11.11 | 0.05 | 42.84 | 40.03 | 3.30 | 37.17 | 0.0 | 2.23 | 100 | ||
| Defects4J | ZSL | 57.01 | 4.05 | 0.06 | 15.72 | 24.30 | 7.50 | 1.09 | 0.0 | 0.91 | 26.90 | 0.79 | 48.55 | 60.70 | 9.92 | 19.65 | 0.0 | 1.09 | 99.64 | |
| FSL | 73.15 | 4.90 | 0.24 | 13.18 | 24.68 | 3.94 | 0.88 | 0.0 | 2.09 | 31.19 | 1.61 | 64.95 | 57.64 | 17.36 | 15.19 | 0.0 | 0.96 | 100 | ||
| CoT | 40.06 | 2.83 | 0.11 | 34.40 | 16.79 | 2.99 | 0.61 | 0.0 | 0.66 | 17.73 | 0.61 | 36.45 | 46.81 | 5.71 | 41.72 | 0.0 | 0.66 | 99.28 | ||
| ToT | 69.56 | 5.52 | 0.32 | 4.57 | 30.60 | 11.20 | 1.03 | 0.0 | 0.79 | 29.81 | 0.79 | 52.92 | 75.95 | 11.20 | 6.23 | 0.0 | 1.0 | 99.97 | ||
| GToT | 29.64 | 2.06 | 0.0 | 52.39 | 12.67 | 2.03 | 0.90 | 0.0 | 1.11 | 7.97 | 0.24 | 28.42 | 32.22 | 1.45 | 55.05 | 0.0 | 1.0 | 99.97 | ||
| CMD | ZSL | 37.72 | 0.0 | 0.0 | 33.33 | 15.79 | 35.96 | 0.0 | 0.0 | 0.0 | 2.63 | 0.0 | 30.70 | 27.19 | 0.0 | 54.39 | 0.0 | 0.0 | 100 | |
| FSL | 18.45 | 0.97 | 0.0 | 49.51 | 25.24 | 14.56 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 18.45 | 16.50 | 0.0 | 77.67 | 0.0 | 0.0 | 100 | ||
| CoT | 6.4 | 0.0 | 0.0 | 78.80 | 7.20 | 2.80 | 0.0 | 0.0 | 0.40 | 0.0 | 0.0 | 6.40 | 7.20 | 0.0 | 89.20 | 0.0 | 0.0 | 100 | ||
| ToT | 15.04 | 0.0 | 0.0 | 71.68 | 7.96 | 18.58 | 0.0 | 0.0 | 0.0 | 0.88 | 0.0 | 17.70 | 15.04 | 0.88 | 76.99 | 0.0 | 0.0 | 100 | ||
| GToT | 19.86 | 0.0 | 0.0 | 58.90 | 15.07 | 17.81 | 0.0 | 0.0 | 0.68 | 1.37 | 0.0 | 18.49 | 21.23 | 0.0 | 74.66 | 0.0 | 0.0 | 100 | ||
| GPT 4 | CMD | ZSL | 30.43 | 1.45 | 0.0 | 11.59 | 14.49 | 20.29 | 0.0 | 0.0 | 1.45 | 2.90 | 0.0 | 36.23 | 36.23 | 0.0 | 34.78 | 0.0 | 0.0 | 98.55 |
| FSL | 41.18 | 1.18 | 0.0 | 10.59 | 31.76 | 20.0 | 0.0 | 1.18 | 0.0 | 2.35 | 0.0 | 48.24 | 45.88 | 2.35 | 36.47 | 0.0 | 0.0 | 100 | ||
| CoT | 14.56 | 0.0 | 0.0 | 28.16 | 18.45 | 17.48 | 0.0 | 0.0 | 0.0 | 0.97 | 0.0 | 18.45 | 35.92 | 0.0 | 54.37 | 0.0 | 0.0 | 98.06 | ||
| ToT | 16.90 | 0.0 | 0.0 | 33.80 | 16.90 | 15.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 23.94 | 22.54 | 0.0 | 53.52 | 0.0 | 0.0 | 97.18 | ||
| GToT | 13.98 | 0.0 | 0.0 | 49.46 | 17.20 | 19.35 | 0.0 | 0.0 | 1.08 | 0.0 | 0.0 | 16.13 | 21.51 | 0.0 | 69.89 | 1.08 | 0.0 | 97.85 | ||
| Mistral 7B | SF110 | ZSL | 63.45 | 9.66 | 0.0 | 0.0 | 14.71 | 7.98 | 0.42 | 1.26 | 1.26 | 31.51 | 0.84 | 65.13 | 80.25 | 10.08 | 3.36 | 0.42 | 0.42 | 100 |
| FSL | 72.69 | 11.89 | 0.0 | 0.88 | 7.05 | 0.44 | 0.0 | 1.32 | 6.17 | 27.75 | 0.0 | 68.28 | 69.60 | 15.86 | 5.29 | 0.0 | 0.0 | 100 | ||
| CoT | 52.26 | 7.04 | 0.0 | 2.51 | 19.10 | 2.01 | 0.5 | 1.51 | 1.51 | 22.61 | 0.5 | 53.27 | 84.92 | 11.56 | 5.53 | 0.0 | 0.5 | 94.47 | ||
| ToT | 42.59 | 2.65 | 0.0 | 8.47 | 12.70 | 2.38 | 0.26 | 0.26 | 1.59 | 10.58 | 0.0 | 43.12 | 42.59 | 8.99 | 11.11 | 0.0 | 0.26 | 92.59 | ||
| GToT | 21.20 | 0.82 | 0.14 | 13.54 | 6.57 | 0.82 | 0.14 | 0.0 | 0.55 | 3.42 | 0.0 | 27.91 | 16.01 | 2.74 | 21.34 | 0.0 | 0.14 | 85.64 | ||
| Defects4J | ZSL | 73.64 | 14.34 | 0.39 | 0.0 | 17.44 | 1.55 | 0.78 | 0.78 | 1.55 | 13.18 | 1.55 | 48.06 | 58.14 | 32.17 | 2.33 | 0.0 | 0.39 | 100 | |
| FSL | 79.35 | 14.17 | 0.4 | 1.62 | 8.5 | 0.81 | 0.4 | 0.0 | 1.21 | 25.51 | 0.81 | 62.75 | 78.95 | 21.46 | 2.02 | 0.4 | 0.4 | 100 | ||
| CoT | 65.13 | 13.82 | 0.0 | 0.66 | 21.05 | 0.66 | 0.0 | 0.0 | 1.32 | 10.53 | 1.32 | 41.45 | 81.58 | 31.58 | 0.66 | 0.0 | 0.0 | 95.39 | ||
| ToT | 69.84 | 7.54 | 0.4 | 1.59 | 22.62 | 2.38 | 0.0 | 0.0 | 1.59 | 22.22 | 1.59 | 64.29 | 78.97 | 23.41 | 3.17 | 0.0 | 0.0 | 98.41 | ||
| GToT | 27.97 | 2.89 | 0.0 | 12.86 | 13.67 | 0.0 | 0.0 | 0.0 | 0.16 | 4.82 | 0.16 | 28.94 | 30.55 | 4.18 | 16.08 | 0.0 | 0.0 | 84.41 | ||
| CMD | ZSL | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
| FSL | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| CoT | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | ||
| ToT | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| GToT | 16.67 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 16.67 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 66.67 | ||
| Mixtral8x7B | CMD | ZSL | 75 | 0.0 | 0.0 | 0.0 | 25.0 | 75.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 25.0 | 25.0 | 0.0 | 0.0 | 0.0 | 0.0 | 100 |
| FSL | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 33.33 | 0.0 | 0.0 | 66.67 | ||
| CoT | 0.0 | 0.0 | 0.0 | 0.0 | 50 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.0 | 0.0 | 50.0 | 0.0 | 0.0 | 100 | ||
| ToT | 44.44 | 11.11 | 0.0 | 44.44 | 11.11 | 0.0 | 0.0 | 0.0 | 0.0 | 11.11 | 0.0 | 55.56 | 33.33 | 0.0 | 55.56 | 0.0 | 0.0 | 100 | ||
| GToT | 0.0 | 0.0 | 0.0 | 50.0 | 50.0 | 50.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.0 | 0.0 | 100 | 0.0 | 0.0 | 100 | ||
| EvoSuite | SF110 | 50.78 | 0.0 | 0.0 | 6.90 | 93.10 | 0.0 | 4.44 | 0.0 | 0.0 | 17.53 | 0.0 | 54.67 | 76.94 | 0.07 | 6.90 | 0.0 | 3.96 | 100 | |
| Defects4J | 38.24 | 0.0 | 0.0 | 6.90 | 93.10 | 0.0 | 4.44 | 0.0 | 0.0 | 10.71 | 0.0 | 62.95 | 82.18 | 6.73 | 0.38 | 0.0 | 4.34 | 100 | ||
@@
∗ AR: Assertion Roulette, CLT: Conditional Logic Test, CI: Constructor Initialization, EM: Empty Test, EH: Exception Handling, GF: General Fixture, MG: Mistery Guest, RP: Redundant Print, RA: Redundant Assertion, SE: Sensitive Equality, ST: Sleepy Test, EA: Eager Test, LT: Lazy Test, DA: Duplicate Assert, UT: Unknown Test, IT: Ignored Test, RO: Resource Optimism, MNT: Magic Number Test, ZSL: Zero-shot Learning, FSL: Few-shot Learning, CoT: Chain-of-Thought, ToT: Tree-of-Thoughts, GToT: Guided Tree-of-Thoughts.